# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v1

@@ Abstract
Key words: Forecast, Data-rich Environment, Hypothesis testing, Shrinkage, Time series, Unit root
@@ 1 Introduction
Prediction is one of the fundamental tasks in statistical analysis of economic and financial data. About one century ago, quantitative analysis and forecasting services were made available to businesses and the general public (Dominguez et al., , 1988). Irving Fisher (1867-1947) spearheaded this initiative, pioneering data-driven forecasting practices grounded in theoretical foundations. His renowned theories, namely the monetary quantity theory, inflation-deflation theory, and index theory, were supported by statistical evidence. For example, Fisher, (1925) was one of the earliest attempts to understand the source of business cycles in association with the price level (inflation), and Fisher, (1926) was the first to report “a statistical relation between unemployment and price changes”, today better known as the Phillips curve. Statistical inference is crucial for integrating economic theory and empirical data, which has become a tradition upheld by successive generations of researchers. In today’s era of big data, we have unprecedented access to a vast amount of digital information about the economy. Recent advancements in statistical inference have uncovered new empirical patterns in prediction practices using datasets with temporal features.
This paper aims at a plain quest: in a high dimensional linear predictive regression model when the number of potential regressors is larger than the sample size, how to conduct asymptotically valid statistical inference for a regressor of particular interest. To the best of our knowledge, no paper has solved this question before. The challenges are twofold. Firstly, conventional predictive regressions were studied in the low-dimensional context. The notable feature of the predictive regression includes the persistence of the regressors (Stambaugh, , 1999), which upends the standard inferential theory based on the standard $t$-statistics. Secondly, to cope with the high dimensionality, one must estimate the coefficient via some regularization method, for example, when the underlying true regression model is sparse, LASSO (Tibshirani, , 1996) is the off-the-shelf method. It is well-known that the LASSO estimator is biased toward zero due to shrinkage, and follows a non-standard asymptotic distribution that is distinct from the asymptotic normal distribution. If we intend to provide an asymptotically normally distributed estimator to facilitate standard statistical inference based on the $t$-statistic, we must simultaneously combat two evils: the Stambaugh bias due to persistent regressors, and the shrinkage bias due to LASSO’s absolute-value penalty.
@@
Literature review. This paper builds on several strands of literature. First, LASSO is one of the most studied methods in recent years, with well-developed theory in high dimension (Bickel et al., , 2009). It is well received and used for economic applications; Belloni et al., (2012), Shi, (2016), Caner and Kock, (2018), and Babii et al., (2022), to name a few. In recent years, the properties of LASSO estimators are studied in nonstationary time series models (Koo et al., , 2020; Lee et al., , 2022; Mei and Shi, , 2024).
Hypothesis testing after LASSO is challenging because of the shrinkage bias. To validate hypothesis testing in high dimensions, Zhang and Zhang, (2014), van de Geer et al., (2014), and Javanmard and Montanari, (2014) have developed the desparsified (debiased) LASSO estimators under the independently and identically distributed (i.i.d.) setting. Adamek et al., (2023) generalize the Dlasso to high dimensional stationary time series inference. We follow this line of desparsified LASSO literature considering its convenience, typically only requiring a baseline regression and an auxiliary regression. An alternative method for statistical inference is Chernozhukov et al., (2018)’s double machine learning (DML), which has been widely used in cross sectional data (Chernozhukov et al., 2022a, ; Chernozhukov et al., 2022b, ; Fan et al., , 2023). DML adopts sample splitting or cross fitting to correct the bias in the initial estimator. In our environment with nonstationary time series, however, it is not clear how to split the sample over the time dimension, due to the time varying nature of the data generating processes.
@@ 2.1 High Dimensional Predictive Regression
The linear model serves as the foundation for understanding basic statistical properties such as consistency and asymptotic distribution, which help applied researchers interpret and communicate their findings effectively. Suppose that a time series of outcome $y_{t}$ is generated by the following linear predictive regression:
@@
where the error term $u_{t}$ is a stationary and martingale difference sequence (m.d.s.) with mean zero and conditional variance $\sigma_{u}^{2}$. [^1] For simplicity of exposition, an intercept in (2.1) is omitted, without loss of generality, as explained by Mei and Shi, (2024) that in LASSO the intercept can be handled by the well-known Frisch-Waugh-Lovell theorem. In practical implementation — throughout all simulations and empirical exercises in this paper — we keep an unpenalized intercept in the model. We consider two types of regressors with different stochastic properties: $X_{t}=(x_{1,t},\cdots,x_{p_{x},t})^{\top}$ is a $p_{x}\times 1$ vector of unit roots satisfying $X_{t}=X_{t-1}+e_{t}$ with $e_{t}=(e_{1,t},\cdots,e_{p_{x},t})^{\top}$ being a $p$-dimensional vector of stationary time series, and $Z_{t}=(z_{1,t},\cdots,z_{p_{z},t})^{\top}$ is a $p_{z}\times 1$ vector of stationary regressors. The two types of regressors are combined into a long vector $W_{t}=(X_{t}^{\top},Z_{t}^{\top})^{\top}=(w_{1,t},\cdots,w_{p,t})^{\top}$ of $p$ ($=p_{z}+p_{x}$) elements, and the associated coefficients are placed into $\theta^{*}=(\beta^{*\top},\gamma^{*\top})^{\top}\in\mathbb{R}^{p}$. For simplicity, let the initial value $\|W_{t=0}\|_{\infty}=O_{p}(1)$.
@@
$\widehat{\theta}^{{\rm S}}:=\arg\min_{\theta}\dfrac{1}{n}\sum_{t=1}^{n}(y_{t}-W_{t-1}^{\top}\theta)^{2}+\lambda\|D\theta\|_{1}.$ (2.2)
@@ 2.2 Two Types of Biases
Following Zhang and Zhang, (2014), we replace the unknown parameter $\theta_{-j}^{*}$ by the feasible workhorse estimator $\widehat{\theta}_{-j}^{{\rm S}}$:
@@
Note that $|{\rm\widehat{B}}_{j}-{\rm B}_{j}|\leq\|g_{j}\|_{\infty}\cdot\|\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}\|_{1}.$ Slasso’s $L_{1}$ estimation error $\|\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}\|_{1}$ is invariant with the choice of score vector $r_{j,t}$, and diminishes if $\widehat{\theta}^{{\rm S}}$ is consistent. Thus, it suffices to control the order of $\|g_{j}\|_{\infty}$ to achieve (R2). The explicit expression of $g_{j}$ in (2.6) suggests that a weak correlation in sup-norm between all $W_{-j,t}$ and the score $r_{j,t}$ (relative to the correlation between $w_{j,t}$ and $r_{j,t}$) helps.
With these routes in mind, we devise XDlasso in the following section. It proceeds with two key steps: (i) conducting an IVX transformation to get a new variable less persistent than the unit root regressors to eliminate the Stambaugh bias; (ii) running an auxiliary LASSO regression to construct an $r_{j,t}$ to cancel out the shrinkage bias.
@@ 2.3 IVX-Desparsified LASSO
For low dimensional predictive regressions, the IVX literature has established asymptotic normality of ${\rm N}_{j}$ in (2.5) taking $r_{j,t}=\widetilde{\zeta}_{j,t}$; see Phillips and Magdalinos, (2009). Furthermore, recall that the vector of regressors $W_{-j,t}$ includes either unit root or stationary regressors. Due to different degrees of persistence, the mildly integrated IV $\widetilde{\zeta}_{j,t}$ and the regressors $W_{-j,t}$ are asymptotically uncorrelated. Thus, when $w_{j,t}$ has a unit root, we can choose the score vector as $r_{j,t}=\widetilde{\zeta}_{j,t}$ to deliver a small-order $\|g_{j}\|_{\infty}$ where $g_{j}$ is defined in (2.6). If $w_{j,t}$ is stationary, however, this score vector fails. When $n$ is large, the instrumental variable $\zeta_{j,t}$ behaves similarly as the stationary regressor $w_{j,t}$, and thus its correlation to the high dimensional stationary regressors in $W_{-j,t}$ is not negligible.
The above analysis implies that we must decorrelate the score vector with the other regressors $W_{-j,t}$ to reduce the order of $g_{j}$ to control for approximation error $\widehat{{\rm B}}_{j}-{\rm B}_{j}$. This decorrelation will produce a unified testing approach for both unit roots and stationary regressors. To this end, we construct a “residual” score vector $\widehat{r}_{j}=(\widehat{r}_{j,0},\cdots,\widehat{r}_{j,n-1})^{\top}$ by the following auxiliary LASSO regression
@@
$\widehat{\varphi}^{(j)}$ $=\arg\min_{\varphi}\dfrac{1}{n}\sum_{t=1}^{n}(\widetilde{\zeta}_{j,t-1}-W_{-j,t-1}^{\top}\varphi)^{2}+\mu\|D_{-j}\varphi\|_{1}$ (2.10)
where $\mu$ is the LASSO tuning parameter. In low dimensional multivariate predictive regressions, IVX transforms each regressor into a less persistent instrumental variable parallel to (2.7), and constructs a two-stage least squares estimator using all the self-generated instrumental variables. In contrast, we only transform the variable of interest $w_{j,t}$ and estimate one auxiliary regression (2.10).
The score vector $\widehat{r}_{j}$ by (2.9) accommodates stationary and unit root regressors. When $w_{j,t}$ is stationary, the magnitude of the instrument $\widetilde{\zeta}{}_{j,t}$ is the same as that of $w_{j,t}$. Thus, the score $\widehat{r}_{j}$ is asymptotically equivalent to the residual of the LASSO regression of $w_{j,t}$ on $W_{-j,t}$. The latter is exactly the score in Zhang and Zhang, (2014) for cross-sectional data, which is also used in Adamek et al., (2023) for stationary time series. When $w_{j,t}$ has a unit root, the instrument $\zeta_{j,t}$ is mildly integrated and has a different degree of persistence from $W_{-j,t}$. Therefore, $\zeta_{j,t}$ is asymptotically uncorrelated to each regressor in $W_{-j,t}$. For notational conciseness, define $\widetilde{W}_{-j,t}=(D_{-j})^{-1}W_{-j,t}$ and $\widetilde{\varphi}^{(j)}=D_{-j}\widehat{\varphi}^{(j)}$. The analysis above suggests $\widetilde{\varphi}^{(j)}\approx 0$ so that $\widehat{r}_{j,t}=\widetilde{\zeta}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\varphi}^{(j)}\approx\widetilde{\zeta}_{j,t}$. Recall from the discussion above (2.8) that the standardized instrument $\widetilde{\zeta}_{j,t}$ is a valid score process for a unit root regressor, and thus the score $\widehat{r}_{j,t}$ can also remove the biases asymptotically. As a result, the residual of the auxiliary LASSO regression provides a unified construction of the score for both stationary and nonstationary $w_{j,t}$. It allows practitioners to conduct hypothesis testing on the coefficient in high dimensional predictive regression regardless of the order of integration of $w_{j,t}$. We maintain an agnostic attitude about the persistence of the regressors; in practical implementation, we need not distinguish these two types.
@@
Step1 Obtain $\widehat{\theta}^{{\rm S}}$ from the Slasso regression (2.2). Save the residual $\widehat{u}_{t}=y_{t}-W_{t-1}^{\top}\widehat{\theta}^{{\rm S}}$ and $\widehat{\sigma}_{u}^{2}=n^{-1}\sum_{t=1}^{n}\widehat{u}_{t}^{2}$. Step2 Obtain the IV $\zeta_{j,t}$ by the transformation (2.7), and standardize it by (2.8). Step3 Run the auxiliary LASSO regression (2.10), and save the residual $\widehat{r}_{j,t}$ in (2.9). Step4 Compute the XDlasso estimator (2.11) and the standard error (2.12). Step5 Obtain the $t$-statistic $t_{j}^{{\rm XD}}=(\widehat{\theta}_{j}^{{\rm XD}}-\theta_{0,j})\big{/}\widehat{\omega}_{j}^{{\rm XD}}.$ (2.13) Reject $\mathbb{H}_{0}$ under the significance level $(1-\alpha)$ if $|t_{j}^{{\rm XD}}|>\Phi_{1-\alpha/2}$, where $\Phi_{a}$ is the $100a$-th percentile of the standard normal distribution.

### v2

@@ Abstract
Key words: Forecast, Data-rich Environment, Hypothesis testing, Shrinkage, Unit root
@@
If modern techniques would have been useless in predicting the last depression, is there any reason to believe they would do better predicting the next? The authors say that the paucity of data available… made it impractical for them to construct a more complicated and more insightful model of the economy. A richer model, Mr. [Ray] Fair concedes, might have yielded better forecasts.”
“Economic Scene; Forecasters’ Art In 1929 and Now” by Peter Passell. New York Times, August 31, 1988,
@@ 1 Introduction
Prediction is one of the fundamental tasks in statistical analysis of economic and financial data. About one century ago, quantitative analysis and forecasting services were made available to businesses and the general public (Dominguez et al.,, 1988). Irving Fisher (1867-1947) spearheaded this initiative, pioneering data-driven forecasting practices grounded in theoretical foundations. His renowned theories, namely the monetary quantity theory, inflation-deflation theory, and index theory, were supported by statistical evidence. For example, Fisher, (1925) was one of the earliest attempts to understand the source of business cycles in association with the price level (inflation), and Fisher, (1926) was the first to report “a statistical relation between unemployment and price changes”, today better known as the Phillips curve. Inference is crucial for integrating economic theory and empirical data, which has become a tradition upheld by successive generations of researchers. In today’s era of big data, we have unprecedented access to a vast amount of digital information about the economy. Recent advancements in inference have uncovered new empirical patterns in prediction practices using datasets with temporal features.
This paper aims at a plain quest: in a high dimensional linear predictive regression model when the number of potential regressors is larger than the sample size, how to conduct asymptotically valid inference for a regressor of particular interest. No paper has solved this question before. The challenges are twofold. Firstly, conventional predictive regressions were studied in the low-dimensional context. The notable feature of the predictive regression includes the persistence of the regressors (Stambaugh,, 1999), which upends the standard inferential theory based on the standard $t$-statistics. Secondly, to cope with the high dimensionality, one must estimate the coefficient via some regularization method, for example, when the underlying true regression model is sparse, LASSO (Tibshirani,, 1996) is the off-the-shelf method. It is well-known that the LASSO estimator is biased toward zero due to shrinkage, and follows a non-standard asymptotic distribution that is distinct from the asymptotic normal distribution. If we intend to provide an asymptotically normally distributed estimator to facilitate standard statistical inference based on the $t$-statistic, we must simultaneously combat two evils: the Stambaugh bias due to persistent regressors, and the shrinkage bias due to LASSO’s absolute-value penalty.
@@
Literature review. This paper builds on several strands of literature. First, LASSO is one of the most studied methods in recent years, with well-developed theory in high dimension (Bickel et al.,, 2009). It is well received and used for economic applications; Belloni et al., (2012), Shi, (2016), Caner and Kock, (2018), and Babii et al., (2022), to name a few. In recent years, the properties of LASSO estimators are studied in various topics in high dimensional time series, including nonstationary time series models (Koo et al.,, 2020; Lee et al.,, 2022; Mei and Shi,, 2024) and inference based on the heteroskedasticity and autocorrelation consistent (HAC) estimation (Babii et al.,, 2020, 2024).
Hypothesis testing after LASSO is challenging because of the shrinkage bias. To validate hypothesis testing in high dimensions, Zhang and Zhang, (2014), van de Geer et al., (2014), and Javanmard and Montanari, (2014) have developed the desparsified (debiased) LASSO estimators under the independently and identically distributed (i.i.d.) setting. Adamek et al., (2023) generalize the Dlasso inference to high dimensional stationary time series. We follow this line of desparsified LASSO literature considering its convenience, typically only requiring a baseline regression and an auxiliary regression. An alternative method for statistical inference is Chernozhukov et al., (2018)’s double machine learning (DML), which has been widely used in cross sectional data (Chernozhukov et al., 2022a, ; Chernozhukov et al., 2022b, ; Fan et al.,, 2024). DML adopts sample splitting or cross fitting to correct the bias in the initial estimator. In our environment with nonstationary time series, however, it is not clear how to split the sample over the time dimension, due to the time varying nature of the data generating processes.
@@ 2.1 High Dimensional Predictive Regression
Suppose that a time series of outcome $y_{t}$ is generated by the following linear predictive regression: [^1] For simplicity of exposition, an intercept in (2.1) is omitted, without loss of generality, as explained by Mei and Shi, (2024) that in LASSO the intercept can be handled by the well-known Frisch-Waugh-Lovell theorem. In practical implementation — throughout all simulations and empirical exercises in this paper — we keep an unpenalized intercept in the model.
@@
where the error term $u_{t}$ is a stationary and martingale difference sequence (m.d.s.) with mean zero and conditional variance $\sigma_{u}^{2}$. We consider two types of regressors with different stochastic properties: $X_{t}=(x_{1,t},\cdots,x_{p_{x},t})^{\top}$ is a $p_{x}\times 1$ vector of unit roots satisfying $X_{t}=X_{t-1}+e_{t}$ with $e_{t}=(e_{1,t},\cdots,e_{p_{x},t})^{\top}$ being a $p$-dimensional vector of stationary time series, and $Z_{t}=(z_{1,t},\cdots,z_{p_{z},t})^{\top}$ is a $p_{z}\times 1$ vector of stationary regressors. The two types of regressors are combined into a long vector $W_{t}=(X_{t}^{\top},Z_{t}^{\top})^{\top}=(w_{1,t},\cdots,w_{p,t})^{\top}$ of $p$ ($=p_{x}+p_{z}$) elements, and the associated coefficients are placed into $\theta^{*}=(\beta^{*\top},\gamma^{*\top})^{\top}\in\mathbb{R}^{p}$. For simplicity, let the initial value $\|W_{t=0}\|_{\infty}=O_{p}(1)$. Define the sparsity index $s=\sum_{j=1}^{p}\boldsymbol{1}\left\{\theta_{j}^{*}\neq 0\right\}$ as the number of nonzero components in the vector of coefficients $\theta^{*}$.
@@
$\widehat{\theta}^{{\rm S}}:=\arg\min_{\theta\in\mathbb{R}^{p}}\dfrac{1}{n}\sum_{t=1}^{n}(y_{t}-W_{t-1}^{\top}\theta)^{2}+\lambda\|D\theta\|_{1}.$ (2.2)
@@ 2.2 Two Types of Biases
Following Zhang and Zhang, (2014), we replace the unknown parameter $\theta_{-j}^{*}$ by the feasible workhorse estimator $\widehat{\theta}_{-j}^{{\rm S}}=(\widehat{\theta}_{k}^{{\rm S}})_{k\neq j}$:
@@
Note that $|{\rm\widehat{B}}_{j}-{\rm B}_{j}|\leq\|g_{j}\|_{\infty}\cdot\|\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}\|_{1}.$ Slasso’s $L_{1}$ estimation error $\|\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}\|_{1}$ is invariant with the choice of score vector $r_{j,t}$, and diminishes if $\widehat{\theta}^{{\rm S}}$ is consistent. Thus, it suffices to control the order of $\|g_{j}\|_{\infty}$ to achieve (R2). The explicit expression of $g_{j}$ in (2.6) suggests that a weak correlation in sup-norm between $W_{-j,t}$ and the score $r_{j,t}$ (relative to the correlation between $w_{j,t}$ and $r_{j,t}$) helps.
With these routes in mind, we devise XDlasso in the following section. It proceeds with two key steps: (i) conducting an IVX transformation to get a new variable less persistent than the unit root regressors to eliminate the Stambaugh bias; (ii) running an auxiliary LASSO regression to construct an $r_{j,t}$ to remove the shrinkage bias.
@@ 2.3 IVX-Desparsified LASSO
For low dimensional predictive regressions, the IVX literature has established asymptotic normality of ${\rm N}_{j}$ in (2.5) taking $r_{j,t}=\widetilde{\zeta}_{j,t}$; see Phillips and Magdalinos, (2009). When $w_{j,t}$ is a unit root, the IV $\zeta_{j,t}$ is mildly integrated and less persistent than a unit root. Furthermore, recall that the vector of regressors $W_{-j,t}$ includes either unit roots or stationary regressors. Due to different degrees of persistence, the mildly integrated IV $\widetilde{\zeta}_{j,t}$ and the regressors $W_{-j,t}$ are asymptotically uncorrelated. Thus, when $w_{j,t}$ has a unit root, we can choose the score vector as $r_{j,t}=\widetilde{\zeta}_{j,t}$ to deliver a small-order $\|g_{j}\|_{\infty}$ where $g_{j}$ is defined in (2.6). If $w_{j,t}$ is stationary, however, this score vector fails. When $n$ is large, the instrumental variable $\zeta_{j,t}$ behaves similarly as the stationary regressor $w_{j,t}$, and thus its correlation to the high dimensional stationary regressors in $W_{-j,t}$ is not negligible.
The above analysis implies that we must decorrelate the score vector with the other regressors $W_{-j,t}$ to reduce the order of $g_{j}$ to control for the approximation error $\widehat{{\rm B}}_{j}-{\rm B}_{j}$. This decorrelation will produce a unified testing approach for both unit roots and stationary regressors. To this end, we construct a “residual” score vector $\widehat{r}_{j}=(\widehat{r}_{j,0},\cdots,\widehat{r}_{j,n-1})^{\top}$ by the following auxiliary LASSO regression
@@
$\widehat{\varphi}^{(j)}$ $=\arg\min_{\varphi\in\mathbb{R}^{p-1}}\dfrac{1}{n}\sum_{t=1}^{n}(\widetilde{\zeta}_{j,t-1}-W_{-j,t-1}^{\top}\varphi)^{2}+\mu\|D_{-j}\varphi\|_{1}$ (2.10)
where $\mu$ is the LASSO tuning parameter, and $D_{-j}={\rm diag}(\{\widehat{\sigma}_{k}\}_{k\neq j})$. In low dimensional multivariate predictive regressions, IVX transforms each regressor into a less persistent instrumental variable parallel to (2.7), and constructs a two-stage least squares estimator using all the self-generated instrumental variables. In contrast, we only transform the variable of interest $w_{j,t}$ and estimate one auxiliary regression (2.10).
The score vector $\widehat{r}_{j}$ by (2.9) accommodates stationary and unit root regressors. When $w_{j,t}$ is stationary, the magnitude of the instrument $\widetilde{\zeta}{}_{j,t}$ shares the same order as that of $w_{j,t}$. Thus, the score $\widehat{r}_{j}$ is asymptotically equivalent to the residual of the LASSO regression of $w_{j,t}$ on $W_{-j,t}$. The latter is exactly the score in Zhang and Zhang, (2014) for cross-sectional data, which is also used in Adamek et al., (2023) for stationary time series. When $w_{j,t}$ has a unit root, the instrument $\zeta_{j,t}$ is mildly integrated and has a different degree of persistence from $W_{-j,t}$. Therefore, $\zeta_{j,t}$ is asymptotically uncorrelated to each regressor in $W_{-j,t}$. For notational conciseness, define $\widetilde{W}_{-j,t}=(D_{-j})^{-1}W_{-j,t}$ and $\widetilde{\varphi}^{(j)}=D_{-j}\widehat{\varphi}^{(j)}$. The analysis above suggests $\widetilde{\varphi}^{(j)}\approx 0$ so that $\widehat{r}_{j,t}=\widetilde{\zeta}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\varphi}^{(j)}\approx\widetilde{\zeta}_{j,t}$. Recall from the discussion below (2.8) that the standardized instrument $\widetilde{\zeta}_{j,t}$ is a valid score process for a unit root regressor, and thus the score $\widehat{r}_{j,t}$ can also remove the biases asymptotically. As a result, the residual of the auxiliary LASSO regression provides a unified construction of the score for both stationary and nonstationary $w_{j,t}$. It allows practitioners to conduct hypothesis testing on the coefficient in high dimensional predictive regression regardless of the order of integration of $w_{j,t}$. We maintain an agnostic attitude about the persistence of the regressors; in practical implementation, we need not distinguish these two types.
@@
Step1 Obtain $\widehat{\theta}^{{\rm S}}$ from the Slasso regression (2.2). Save the residual $\widehat{u}_{t}=y_{t}-W_{t-1}^{\top}\widehat{\theta}^{{\rm S}}$ and $\widehat{\sigma}_{u}^{2}=n^{-1}\sum_{t=1}^{n}\widehat{u}_{t}^{2}$. Step2 Obtain the IV $\zeta_{j,t}$ by the transformation (2.7), and standardize it by (2.8). Step3 Run the auxiliary LASSO regression (2.10), and save the residual $\widehat{r}_{j,t}$ in (2.9). Step4 Compute the XDlasso estimator (2.11) and the standard error (2.12). Step5 Obtain the $t$-statistic $t_{j}^{{\rm XD}}=(\widehat{\theta}_{j}^{{\rm XD}}-\theta_{0,j})\big{/}\widehat{\omega}_{j}^{{\rm XD}}.$ (2.13) Reject $\mathbb{H}_{0}$ under the significance level $\alpha$ if $|t_{j}^{{\rm XD}}|>\Phi_{1-\alpha/2}$, where $\Phi_{a}$ is the $100a$-th percentile of the standard normal distribution.